# Understanding the approach to thermalization in a non-Abelian gauge theory

### v1

@@ 3.3 Chaotic dynamics near deconfinement phase transition
Near the deconfinement phase transition temperature $T_{c}$ the hard, electric and the magnetic scales coincide, hence the effective thermal field theory described in the preceding section can no longer describe the gauge field dynamics. Hence we change our strategy to estimate the Lyapunov exponent. Since SU(2) gauge theory shares the same universality class as the 3D Ising model, hence its long-wavelength or low energy critical modes modes will undergo the same dynamics as a $Z_{2}$ scalar field. We thus simulate a relativistic $Z_{2}$ scalar field theory to understand the chaotic dynamics of the critical modes near the deconfinement phase transition. We start with the Hamiltonian which on the lattice is defined as,
@@
Here $\sum_{\mathbf{y\sim x}}$ implies the sum over all nearest neighbors $\mathbf{y}$ around the lattice site $\mathbf{x}$. In our work, we set the values of $m^{2}a^{2}=-1~{},~{}\lambda_{s}=1$. We first generate thermal ensembles using the Langevin algorithm, in which we consider the time evolution of the scalar and its conjugate momentum fields through equations of motion,
@@
where $\gamma$ is the damping term and $\eta_{\mathbf{x}}$ are Gaussian noise fields defined on each lattice site with a zero mean value which satisfies $\langle\eta_{\mathbf{x}}(t)\eta_{\mathbf{y}}(t^{\prime})\rangle=\delta_{\mathbf{xy}}\delta(t-t^{\prime})$. The relation between the damping factor $\gamma$ and the coefficient of the noise term due to thermal fluctuation can be obtained using the fluctuation-dissipation theorem. To initialize the algorithm we select initial values of $\phi_{\mathbf{x}}(0)$ and $\pi_{\mathbf{x}}(0)$ on each lattice site $\mathbf{x}$. We then numerically evolve the above differential equations using the symplectic leap-frog algorithm with a time step $\Delta t=0.01$. After some Langevin time $t_{L}$, we expect the system to thermalize since the Langevin algorithm is ergodic and follows the condition of detailed balance. Once we achieve thermalization using the Langevin algorithm, we use the thermal configuration thus obtained, switch off the noise and damping terms and perform a real-time classical evolution of the scalar fields. We also additionally select infinitesimal perturbation about the initial thermal configurations denoted by $\delta\phi_{\mathbf{x}}(0)=0$ and $\delta\pi_{\mathbf{x}}(0)=r\delta_{\mathbf{x0}}$, where $r$ is a Gaussian random number between 00-$0.01$. Since $\delta\pi_{\mathbf{x}}(t)=\partial_{t}\delta\phi_{\mathbf{x}}(t)$ one can re-express the initial momentum profile in the configuration space at $\mathbf{x}\equiv(0,0,0),t=0$ and then evolve the fluctuations of the scalar field through
@@
In the evolution equation Eq. 11 one requires the values of the scalar fields $\phi_{\mathbf{x}}(t)$ which are obtained through a classical Hamiltonian evolution of initial thermalized values of $\phi_{\mathbf{x}}(0)$ at each temperature $T$, generated earlier using the Langevin algorithm. We further evolve the initial momentum field fluctuations $\delta\pi_{\mathbf{x}}(t)$ through $\partial_{t}\delta\phi_{\mathbf{x}}(t)=\delta\pi_{\mathbf{x}}(t)$. We thus have all ingredients to calculate the OTOC function in semiclassical limit using Eq. 2, from which we extract a positive Lyapunov exponent, discussed in the next section.
Figure 1: The Lyapunov exponent as a function of $T/T_{c}$ for a $Z_{2}$ scalar field theory which from universality represent the critical long-distance modes of a $SU(2)$ thermal state near the deconfinement temperature $T_{c}$ The inset shows the growth of OTOC functions with time $t.T_{c}$.
@@ 4 Results
We first measure the OTOC functions for $Z_{2}$ scalar field theory at different temperatures close to the phase transition, which is shown in the inset of Fig 1. After an initial dip, The OTOCs exhibit an exponential rise as a function of time. By performing a fit to the OTOCs at late times $t.T_{c}>1000$ with an exponential ansatz, we extract the (maximal) Lyapunov exponent at different temperatures, results of which are shown in Figure 1. Here $a.T_{c}=9.37$ in dimensionless unit and agrees very well with the values in the literature [28]. The Lyapunov exponent thus extracted varies as $\lambda_{L}\sim T^{6.8(7)}$ below $T_{c}$, whereas in the range $(1,1.05)T_{c}$ its temperature variation can as characterized as $T^{-2.8(3)}$, which was determined after performing a fit to our data. The deconfinement phase transition temperature in SU(2) is $T_{c}\sim 300$ MeV [27].
@@
Further in the inset of Fig 2 we show the heat-map of the OTOC at $T=T_{c}$ in a two dimensional plane denoted by the time and one of the spatial coordinates, performing a sum over the other spatial coordinates. The OTOC spreads ballistically within the light cone from an initial small perturbation at the origin as evident from the color gradient of the heat-map. In general the $\lambda_{L}$ is a function of the velocity and it can be measured from the OTOC function using Eq. 4. As evident in Fig 1, the Lyapunov exponent is maximum along $v=0$ i.e. along the time direction. We next calculate the butterfly velocity $v_{B}$ by fitting the OTOC $C(\mathbf{x},t)$ to the ansatz given in Eq. 4. The $v_{B}$ varies across the phase transition and is found to be maximum at $T=T_{c}$. In Fig 2, we have shown the diffusion coefficient as a function of temperature, which is related to the butterfly velocity through the relation $D=v_{B}^{2}/\lambda_{L}$ [29]. As evident from the figure, the diffusion coefficient is constant below $T_{c}$ but decreases with increasing temperatures as $(T/T_{c})^{-3.7(2)}$ for $T>T_{c}$.
@@
Figure 2: The temperature dependence of the diffusion coefficient $D$ near $T_{c}$ extracted using Eq. 4. The inset shows the ballistic spread of the OTOC function of $Z_{2}$ scalar fields in $x$-$t$ plane at $T_{c}$. The color profile from green to red denotes increasing values of OTOC in the $x$-$t$ plane.
@@
At higher temperatures, $T\gg T_{c}$, the gauge invariant distance measure $d(t)=\frac{1}{2N_{p}}\sum_{P}|\text{tr}U_{P}(t)-\text{tr}U_{P}^{\prime}(t)|$ were measured on lattice of $N=32$ spatial sites and for different temperatures results for which are shown in the inset of Fig 3. The initial $d(t)\sim 10^{-15}$ measures the typical difference between the plaquettes in two thermal configurations which, when evolved in time grows exponentially eventually saturating at late times to $\sim 10^{-2}$, since the trace of plaquette is bounded to unity and the gauge group is compact. The lattice size are chosen large enough for this purpose such that the values are $\sim 15(6)$ fm at temperatures $0.6(1.5)$ GeV respectively. Extracting the (maximal) Lyapunov exponent $\lambda_{L}$ from the exponentially rising part of $d(t)$ with time, we study its dependence on temperature, results of which are shown in Fig. 3. The $\lambda_{L}$ varies linearly with temperature between $0.6$-$3$ GeV with a slope $\lambda_{L}/T\sim 0.52$ which satisfies the conjectured bound $\lambda_{L}\leq 2\pi T$ in quantum many-body chaotic systems [30] and also satisfies $\lambda_{L}\sim g^{2}\mathcal{E}/6$, where $\mathcal{E}$ is the average energy per plaquette [11]. The diffusion coefficient extracted from $\lambda_{L}$ thus vary inversely with temperature as $(T/T_{c})^{-0.95(3)}$ shown in Fig  4. At these temperatures the distance measure spreads from an initial tiny fluctuation at $t.T=0$ ballistically over the space-time and grows exponentially within the light-cone, evident from the inset of the same figure, the color gradient from green to red indicates increasing values of the distance measure $d(t)$. Across these high temperatures, the butterfly velocity is $v_{B}=0.8c$ and does not vary significantly. Compared to its value at $T_{c}$, the butterfly velocity reduces by only $\sim 10\%$ which implies that once thermalized, most of the configuration (color) space gets occupied.
@@
Figure 3: The Lyapunov exponent as a function of $T/T_{c}$ for a SU(2) gauge theory at high temperatures in thermal equilibrium shown as triangles. These are compared with the Lyapunov exponents in a non-thermal state of SU(2) with comparable energy densities (circles). The inset shows the growth of the gauge invariant distance measure with time $t.T_{c}$ for the thermal SU(2) state for 3 different temperatures.
@@
We next extract the Lyapunov exponent within the self-similar scaling regime of the non-equilibrium SU(2) plasma as outlined in Sec. 3.1. In order to compare with a typical thermal state at temperature $T\sim 2$ GeV, we set the initial density of gluons to be $n_{0}=16$, and consider lattice box of spatial extent $\sim 16$ fm such that the energy densities are similar in the thermal as well in the non-thermal scaling regime. The (maximal) Lyapunov exponent $\lambda_{L}=0.66~{}Q_{s}$ is similar in magnitude to that in a thermal state at $T\gg T_{c}$. We have also shown the values of the maximal Lyapunov exponents in Fig 3 in this non-thermal state, as a function of an effective temperature which is the fourth root of the energy density, as blue data points. The data points lie on the same curve determining the temperature dependence of the thermal Lyapunov exponents. For lower energy densities $\varepsilon(\sim\sqrt{n_{0}}\lesssim 1)$ it was earlier conjectured [31] and later observed [15] from classical-statistical lattice calculations in the self-similar non-thermal scaling regime of SU(3), that the (maximal) Lyapunov exponent $\lambda_{L}$ varies as $\varepsilon^{1/4}$. Our calculations were performed in SU(2) non-thermal scaling state at larger gluon energy densities $(\sim\sqrt{n_{0}}\gtrsim 4)$ where we observe a linear scaling of $\lambda_{L}$ with $\varepsilon$. The linear scaling of $\lambda_{L}$ at higher energies similar to a thermal plasma can be understood in terms of linear temperature dependence of the plasmon damping rate [32].
Figure 4: The temperature dependence of the diffusion coefficient $D$ in SU(2) gauge theory at high temperatures $T\gg T_{c}$. The inset shows the ballistic spread of the Lyapunov exponents in a typical thermal state of SU(2) at $T=1.2$ GeV. The color gradient from green to red denotes increasing values of OTOC in the $x$-$t$ plane.
@@ 5 Physical implications of our results
Our results demonstrate that the infrared (soft) modes of the SU(2) have chaotic dynamics both in thermal equilibrium at a wide range of temperatures $0.9$-$10~{}T_{c}$ as well in a particular non-thermal fixed point as evident from the positive (maximal) Lyapunov exponent $\lambda_{L}$. Interestingly the $\lambda_{L}$ decreases near critical temperature $T_{c}$ with temperatures $T>T_{c}$, maximizing at the deconfinement phase transition. Such a property of $\lambda_{L}$ for the critical modes has been observed earlier in spin systems  [33, 34, 35, 36] as well as in scalar field theories in 2 dimensions [18]. The temperature dependence of $\lambda_{L}$ thus, is not particularly related to the details of microscopic interactions between the degrees of freedom but to the symmetries of the system.
The diffusion coefficient $D$ which measures how fast the correlation between the fields increases with time, has many interesting features as a function of temperature. The $\lambda_{L}$ at $T_{c}$ shown in Fig. 1 correspond to the long-wavelength critical modes of SU(2) gauge theory which are a subset of the infrared magnetic modes, whose momenta are bounded by $|\mathbf{p}|\leq g^{2}T/\pi$. The fact that $D$ is independent of temperature below and at $T_{c}$ and decreases sharply above $T_{c}$ implies that these critical gluon modes at the phase transition are maximally spread in the configuration space. This spread is an inherent chaotic phenomena, which also manifests itself in the stochastic properties [37] of the eigenvalue density of Wilson loop operator in the limit of large number of colors. In this limit, the quantum fluctuations are sub-dominant, and an order-disorder transition [38, 39] in terms of the area of Wilson loops has similar features as the (de)confinement transition.
The temperature dependence of $D$ at $T\gg T_{c}$ in Fig. 4 can be interpreted from the fact that $D$ represents a diffusion in the configuration space of color gauge links and hence can be related to the inverse of color conductivity $\sigma$. The $\sigma$ calculated perturbatively in a thermal non-Abelian plasma, increases linearly with temperature [26]. In the effective dynamics of the soft (magnetic) modes given in Eq. 7, the $\sigma$ represent how these modes rearrange in response to the ultra-relativistic hard modes. The soft modes receive random kicks from the thermal bath of hard modes and a finite $\sigma$ effectively dampens them, allowing for an efficient thermalization.
@@
for the range of temperatures $2$-$10~{}T_{c}$, considering that about $1/3$ of $18N^{3}$ number of the Lyapunov exponents are positive [12] and their values can be obtained from the relation $\lambda(v)\sim\lambda_{L}[1-(v/v_{B})^{2}]$ evident in our data, where we have taken the values of $v$ as randomly distributed between 00 and $v_{B}$. We have verified that our estimate does not vary significantly if the values of $v$ are chosen to be uniformly distributed between $[0,v_{B}]$. Starting from a thermal state $\dot{s}_{KS}$ measures the rate at which information about the initial state of the system is lost. We use the thermodynamic relation $s_{\text{th}}T=\varepsilon+p$ to obtain the entropy density of a thermal state of SU(2) to be $s_{\text{th}}a^{3}=\frac{4}{3}\frac{\varepsilon.a^{4}}{T.a}$ at temperatures $T\gg T_{c}$ where the energy density and pressure are related by $\varepsilon\simeq 3p$. On the other hand the entropy density for the non-thermal state in the self-similar regime described in section 3.1, $s_{\text{non-th}}a^{3}=-\int{\tilde{f}(|\mathbf{p}|)\ln\left[\tilde{f}(|\mathbf{p}|)\right]~{}d^{3}(a\mathbf{p})}\simeq 4.8$ can be calculated by performing a numerical integration using the non-thermal distribution $f$ obtained in our computations. Using the entropy densities of the thermal and non-thermal states one can derive the thermalization time $t_{\text{th}}=0.8/\lambda_{L}(T)$, by integrating over Eq. 12. Thus starting from a highly occupied gluon state in a non-thermal scaling regime, a thermal state at $T\simeq 600$ MeV can be reached within a time, $t_{\text{th}}\sim 0.50(3)$ fm/c, whereas it would take $t_{\text{th}}\sim 0.70(5)$ fm/c to reach a thermal state at $T\simeq 450$ MeV.

### v2

@@ 3.3 Chaotic dynamics near deconfinement phase transition
Near the deconfinement phase transition temperature $T_{c}$ the hard, electric and magnetic scales coincide, hence the effective thermal field theory described in the preceding section can no longer describe the gauge field dynamics. Hence we change our strategy to estimate the Lyapunov exponent. Since SU(2) gauge theory shares the same universality class as the 3D Ising model [31], hence its long-wavelength or low energy critical modes will undergo the same dynamics as a $Z_{2}$ scalar field. We thus simulate a relativistic $Z_{2}$ scalar field theory to understand the chaotic dynamics of the critical modes near the deconfinement phase transition. We start with the Hamiltonian defined on the lattice as,
@@
Here $\sum_{\mathbf{y\sim x}}$ implies the sum over all nearest neighbors $\mathbf{y}$ around the lattice site $\mathbf{x}$. In our work, we set the values of $m^{2}a^{2}=-1~{},~{}\lambda_{s}=1$. We first generate thermal ensembles using the Langevin dynamics, in which we consider the time evolution of the scalar and its conjugate momentum fields through equations of motion,
@@
where $\gamma$ is the damping term and $\eta_{\mathbf{x}}$ are Gaussian noise fields defined on each lattice site with a zero mean value and satisfying $\langle\eta_{\mathbf{x}}(t)\eta_{\mathbf{y}}(t^{\prime})\rangle=\delta_{\mathbf{xy}}\delta(t-t^{\prime})$. The relation between the damping factor $\gamma$ and the coefficient of the noise term mimicking thermal fluctuations can be obtained using the fluctuation-dissipation theorem. To initialize the algorithm, we select random initial values of $\phi_{\mathbf{x}}(0)$ and $\pi_{\mathbf{x}}(0)$ on each lattice site $\mathbf{x}$. We then numerically evolve the above differential equations using the symplectic leap-frog integrator with a time step $\Delta t=0.01$. Since Langevin dynamics is ergodic and follows the condition of detailed balance, we expect the system to thermalize after some Langevin time $t_{L}$. Once we achieve thermalization using the Langevin dynamics, we take the thermal configurations of the scalar fields and perform a real-time classical evolution. Further, we implement infinitesimal perturbation about the initial thermal configurations, denoted by $\delta\phi_{\mathbf{x}}(0)=0$ and $\delta\pi_{\mathbf{x}}(0)=r\delta_{\mathbf{x0}}$, where $r$ is a Gaussian random number between 00 and $0.01$. Since $\delta\pi_{\mathbf{x}}(t)=\partial_{t}\delta\phi_{\mathbf{x}}(t)$ one can re-express the initial momentum profile in the configuration space at $\mathbf{x}\equiv(0,0,0)$ and $t=0$ and then evolve the fluctuations of the scalar field via
@@
In Eq. 14 one requires the values of the scalar fields $\phi_{\mathbf{x}}(t)$ as an input which are obtained through a classical Hamiltonian evolution of initial thermalized values of $\phi_{\mathbf{x}}(0)$ at each temperature $T$. We further evolve the initial momentum field fluctuations $\delta\pi_{\mathbf{x}}(t)$ using $\partial_{t}\delta\phi_{\mathbf{x}}(t)=\delta\pi_{\mathbf{x}}(t)$. We thus have all the ingredients to calculate OTOC in the semiclassical limit using Eq. 2, from which we extract the Lyapunov exponent, as discussed in the next section.
Figure 1: The Lyapunov exponent as a function of $T/T_{c}$ for a $Z_{2}$ scalar field theory, which from universality represent the critical long-distance modes of a $SU(2)$ thermal state near the deconfinement transition temperature $T_{c}$. The inset shows the growth of OTOC with time $t.T_{c}$.
@@ 4 Results
We first measure the OTOCs for $Z_{2}$ scalar field theory at different temperatures close to the phase transition, shown in the inset of Fig 1. After an initial dip, the OTOCs exhibit an exponential rise as a function of time. By performing a fit to the OTOCs at late times $t.T_{c}>1000$ with an exponential ansatz, we extract the (maximal) Lyapunov exponent at different temperatures, results of which are shown in Figure 1. The Lyapunov exponent thus extracted varies as $\lambda_{L}\sim T^{6.8(7)}$ below $T_{c}$, whereas in the range $(1,1.05)T_{c}$ its temperature variation can as characterized as $T^{-2.8(3)}$, which was determined after performing a fit to our data. In our calculations $a.T_{c}=9.37$ which agrees very well with the values obtained in the literature [32]. For reference, the deconfinement phase transition temperature in SU(2) is $T_{c}\sim 300$ MeV [30].
@@
Further, in the inset of Fig 2, we show the heat-map of the OTOC at $T=T_{c}$ in a 2-dimensional plane as a function of time $t$ and one of the spatial coordinates, performing a sum over the other spatial coordinates. The OTOC spreads ballistically within the light cone from an initial small perturbation at the origin as evident from the color gradient of the heat-map. In general, $\lambda_{L}$ is a function of the velocity and it can be measured from the OTOC using Eq. 4. As evident in Fig 1, the Lyapunov exponent is maximum along $\mathrm{v}=0$ i.e. along the time direction where the rate of increase in the velocity is the largest compared to any other direction. We next calculate the butterfly velocity $\mathrm{v}_{B}$ by fitting the OTOC $C(\mathbf{x},t)$ to the ansatz given in Eq. 4. The $\mathrm{v}_{B}$ varies across the phase transition and is found to be maximum at $T=T_{c}$. In Fig 2, we have shown the diffusion coefficient as a function of temperature, which is related to the butterfly velocity through the relation in Eq. 5. As evident from the figure, the diffusion coefficient is constant below $T_{c}$ but decreases with increasing temperatures as $(T/T_{c})^{-3.7(2)}$ for $T>T_{c}$.
@@
Figure 2: The temperature dependence of the diffusion coefficient $D$ for $Z_{2}$ scalar fields near $T_{c}$, extracted using Eq. 4. The inset shows the ballistic spread of the OTOC of $Z_{2}$ scalar fields in $x$-$t$ plane at $T_{c}$. The color profile from green to red denotes increasing values of OTOC in the $x$-$t$ plane.
@@
At higher temperatures, $T\gg T_{c}$, the gauge invariant distance measure $d(t)$ was measured on a lattice with $N=32$ spatial sites and for different temperatures, results for which are shown in the inset of Fig 3. The initial $d(t)\sim 10^{-15}$ measures the typical difference between the plaquettes in two thermal configurations which, when evolved in time, grows exponentially eventually saturating at late times to $\sim 10^{-2}$, since the trace of plaquette is bounded to unity and the gauge group is compact. Within this timescale, the momentum mode occupancies of gluons remain similarly close to the thermal distribution. The lattice size is chosen to be large enough such that the physical lengths are $\sim 15(6)$ fm at temperatures $0.6(1.5)$ GeV respectively. We extract the (maximal) Lyapunov exponent $\lambda_{L}$ from the exponentially rising part of $d(t)$ and study its temperature dependence, results of which are shown in Fig. 3. The $\lambda_{L}$ varies linearly with temperature in the range $0.6$-$3$ GeV with a slope $\lambda_{L}/T\sim 0.52$ which satisfies the conjectured bound $\lambda_{L}\leq 2\pi T$ in quantum many-body chaotic systems [33] and also satisfies $\lambda_{L}\sim g^{2}\mathcal{E}/6$, where $\mathcal{E}$ is the average energy per plaquette [11]. The diffusion coefficient extracted from $\lambda_{L}$ using Eq. 5 varies inversely with temperature as $(T/T_{c})^{-0.95(3)}$ shown in Fig  4. At these temperatures the distance measure spreads from an initial tiny fluctuation at $t.T=0$ spreads ballistically over the space-time and grows exponentially within the light-cone, evident from the inset of the same figure, the color gradient from green to red indicates increasing values of the distance measure $d(t)$. Across these high temperatures, the butterfly velocity is $\mathrm{v}_{B}=0.8c$ and does not vary significantly. Compared to its value at $T_{c}$, the butterfly velocity reduces by only $\sim 10\%$ which implies that once thermalized, most of the configuration (color) space gets occupied.
@@
Figure 3: The Lyapunov exponent as a function of $T/T_{c}$ for a SU(2) gauge theory in thermal equilibrium at high temperatures $T>>T_{c}$, shown as triangles. These are compared with the Lyapunov exponents in a non-thermal state of SU(2) with comparable energy densities (circles). The inset shows the growth of the gauge invariant distance measure with time $t.T_{c}$ for a thermal state of SU(2) for 3 different temperatures.
@@
We next extract the Lyapunov exponent within the self-similar scaling regime of the non-equilibrium SU(2) plasma as outlined in Sec. 3.1. In order to compare with a typical thermal state at temperature $T\sim 2$ GeV, we set the initial density of gluons to be $n_{0}=16$, and consider lattice box of spatial extent $\sim 16$ fm such that the energy densities are similar in the thermal as well in the non-thermal scaling regime. The (maximal) Lyapunov exponent $\lambda_{L}=0.66~{}Q_{s}$ extracted in this non-thermal state is similar in magnitude to that in a thermal state at $T\sim 2$ GeV. We have also shown the values of the maximal Lyapunov exponents in Fig 3 in this non-thermal state, as a function of the effective temperature extracted from the fourth root of the energy density, as blue data points. The data points lie on the same curve that characterizes the temperature dependence of the thermal Lyapunov exponents. For lower energy densities $\varepsilon(\sim\sqrt{n_{0}}\lesssim 1)$ it was earlier conjectured [34] and later observed [15] from classical-statistical lattice calculations in the self-similar non-thermal scaling regime of SU(3), that the (maximal) Lyapunov exponent $\lambda_{L}$ varies as $\varepsilon^{1/4}$. Our calculations performed in the non-thermal scaling regime of SU(2), with a larger initial gluon energy density $(\sim\sqrt{n_{0}}\gtrsim 4)$ also yield a linear scaling of $\lambda_{L}$ with $\varepsilon^{1/4}$. The linear scaling of $\lambda_{L}$ at higher energies similar to a thermal plasma can be understood in terms of linear temperature dependence of the plasmon damping rate [35].
@@ 5 Physical implications of our results
Our results demonstrate that the infrared (soft) modes of SU(2) gauge theory exhibit chaotic dynamics both in thermal equilibrium at a wide range of temperatures $0.9$-$10~{}T_{c}$ as well in a particular non-thermal fixed point as evident from the positive (maximal) Lyapunov exponent $\lambda_{L}$. Interestingly the $\lambda_{L}$ decreases beyond the critical temperature $T_{c}$ with temperatures $T>T_{c}$, maximizing at the deconfinement phase transition. Such a property of $\lambda_{L}$ for the critical modes has been observed earlier in spin systems  [36, 37, 38, 39] as well as in scalar field theories in 2 dimensions [19]. The temperature dependence of $\lambda_{L}$, thus, is not particularly related to the details of microscopic interactions between the degrees of freedom but to the symmetries of the system.
Figure 4: The temperature dependence of the diffusion coefficient $D$ in SU(2) gauge theory at high temperatures $T\gg T_{c}$. The inset shows the ballistic spread of the Lyapunov exponents in a typical thermal state of SU(2) at $T=1.2$ GeV. The color gradient from green to red denotes increasing values of the distance function in the $x$-$t$ plane.
The diffusion coefficient $D$ which measures how fast the correlation between the fields increases with time, has many interesting features as a function of temperature. The $\lambda_{L}$ at $T_{c}$ shown in Fig. 1 correspond to the long-wavelength critical modes of SU(2) gauge theory which are a subset of the infrared magnetic modes, whose momenta are bounded by $|\mathbf{p}|\leq g^{2}T/\pi$. The fact that $D$ is independent of temperature below and at $T_{c}$ and decreases sharply above $T_{c}$ implies that these critical gluon modes at the phase transition are maximally spread in the configuration space. This spread is an inherently chaotic phenomena, which also manifests itself in the stochastic properties [40] of the eigenvalue density of Wilson loop operator in the limit of large number of colors. In this limit, the quantum fluctuations are sub-dominant, and an order-disorder transition [41, 42] in terms of the area of Wilson loops has similar features as the (de)confinement transition.
The temperature dependence of $D$ at $T\gg T_{c}$ as shown in Fig. 4 can be interpreted from the fact that $D$ represents a diffusion in the configuration space of color gauge links and hence can be related to the inverse of color conductivity $\sigma$. The color conductivity is related to the color diffusion coefficient through the relation, $\sigma=\chi D$ [43, 44], where $\chi$ is the color charge susceptibility. The $\sigma$ calculated using perturbative techniques in a thermal non-Abelian plasma, increases linearly with temperature [29] and $\chi$ varies as $\sim T^{2}$. Hence the diffusion coefficient varies as $D\sim\frac{1}{T}$, which parametrically depends on the inverse of color conductivity $\sigma$. In the effective dynamics of the soft (magnetic) modes described by Eq. 9, $\sigma$ represent how these modes rearrange in response to the ultra-relativistic hard modes. The soft modes receive random kicks from the thermal bath consisting of hard modes and a finite $\sigma$ effectively dampens them, allowing for efficient thermalization.
@@
for the range of temperatures $2$-$10~{}T_{c}$, considering that about $1/3$ of $18N^{3}$ number of the Lyapunov exponents are positive [12] and their values can be obtained from the relation $\lambda(v)\sim\lambda_{L}[1-(\mathrm{v}/\mathrm{v}_{B})^{2}]$ evident in our data, where we have taken the values of $\mathrm{v}$ as randomly distributed between 00 and $\mathrm{v}_{B}$. We have also verified that our estimate does not vary significantly if the values of $\mathrm{v}$ are chosen to be uniformly distributed between $[0,\mathrm{v}_{B}]$. Starting from a thermal state, $\dot{s}_{KS}$ measures the rate at which information about the initial state of the system is lost. We use the thermodynamic relation $s_{\text{th}}T=\varepsilon+p$ to obtain the entropy density of a thermal state of SU(2) to be $s_{\text{th}}a^{3}=\frac{4}{3}\frac{\varepsilon.a^{4}}{T.a}$ at temperatures $T\gg T_{c}$ where the energy density and pressure are related by $\varepsilon\simeq 3p$. On the other hand the entropy density for the non-thermal state in the self-similar regime described in section 3.1, $s_{\text{non-th}}a^{3}=-\int{\tilde{f}(|\mathbf{p}|)\ln\left[\tilde{f}(|\mathbf{p}|)\right]~{}d^{3}(a\mathbf{p})}\simeq 4.8$ can be calculated by performing a numerical integration using the non-thermal distribution $f$ obtained in our computation. Using the entropy densities of the thermal and non-thermal states one can derive the thermalization time $t_{\text{th}}=0.8/\lambda_{L}(T)$, by integrating over Eq. 15. Thus starting from a highly occupied gluon state in a non-thermal scaling regime, a thermal state at $T\simeq 600$ MeV can be reached within a time, $t_{\text{th}}\sim 0.50(3)$ fm/c, whereas it would take $t_{\text{th}}\sim 0.70(5)$ fm/c to reach a thermal state at $T\simeq 450$ MeV.